OpenStudy (anonymous):

what is the solution to the following equation: x^2+14x+48=0

OpenStudy (anonymous):

which method you would you like to get the solution a) by quadratic formula b) by grouping method c) by method of completing square

OpenStudy (anonymous):

which is easiest?

OpenStudy (anonymous):

-8, -6

OpenStudy (anonymous):

by formula

OpenStudy (anonymous):

ok...well never mind I got it thx

OpenStudy (anonymous):

factor it, (x+8) (x+6) and set both = to zero x+8=0 x+6=0 x=-8 x=-6

OpenStudy (haseeb96):